OpenStudy (anonymous):

Jackson takes 6 hours to clean the garage alone,but can clean it in 1.5 hours if he works with Mary.how long does it take Mary to clean the garage if she works alone

OpenStudy (whpalmer4):

Why did you close and repost?

OpenStudy (anonymous):

ipad went dead now on my computer

OpenStudy (tkhunny):

This does not answer the question. Think in different terms. You ar given 6 hours per job, Ponder on jobs / 6 hrs

OpenStudy (whpalmer4):

Okay, as I said before, Jackson takes 6 hours to do 1 garage, so he does 1/6 of a garage every hour. In 1.5 = 3/2 hours, he does (3/2)*(1/6) = (1/4) of a garage. But working with Mary, in 1.5 = 3/2 hours, they manage to do a whole garage. If Jackson only did 1/4 of the garage (the slacker), how much did Mary do?